What is A if P = [A × (1+R)L] + [B × (1+R)M] and P = 26,670, B = 6,461, R = 0.1008, L = 2, and M = 9? Round answer to nearest whole number.

Please show work
Answer #1

P = [A × (1+R)L] + [B × (1+R)M]

P = 26,670, B = 6,461, R = 0.1008, L = 2, and M = 9

26,670 = A * (1 + 0.1008)^2 + 6,461 * (1 + 0.1008)^9

26,670 = A * 1.21176064 + 6,461 * 2.37342650515

26,670 = A * 1.21176064 + 15,334.7086497742

A * 1.21176064 = 26,670 - 15,334.7086497742

A * 1.21176064 = 11,335.2913502

A = 11,335.2913502/1.21176064

A = 9,354

Can you please upvote? Thank You :-)